# Construction of a Basis of Kähler differentials for generic plane branches

### v1

@@ 6.2.3. $\delta_{y}=1$ and $\beta_{0}=n$
Again, $l_{0}=0$ for $\mathcal{C}$, $E=\{xy=0\}$ and ${E}_{1}=\{x_{1}=0\}$. As described in subsubsection 4.2.2, ${\mathcal{T}}_{0}=(\Omega_{1},\Omega_{2})$ where $\Omega_{1}=xdy\in\Delta_{n,<}^{0}$, $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{\mathcal{C}}(\Omega_{1})=2n$ and $\nu_{\mathcal{C}}(\Omega_{2})=n+\beta_{1}$. We also have $\tilde{\mathcal{T}}_{0}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$ where $\tilde{\Omega}_{1}=dx_{1}\in\Delta_{n,<}$, $\tilde{\Omega}_{2}=x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{1})=n$ and $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2})=\beta_{1}$. Once more, the families $\mathcal{T}_{0}$ and $\tilde{\mathcal{T}}_{0}$ are both terminal, and parallel. Equation (35) for $j\in\{1,2\}$ holds too.
@@ 6.2.4. $\delta_{y}=1$ and $\beta_{0}\not\in(n)$
Take $1\leq k\leq 2{\nu}_{l+1}$, and assume $\Omega_{k}^{1}\neq\Omega_{k}^{0}$. Then there exists $1\leq j<k$ such that $[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}$. Since
$[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}=[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0})]_{n},$
@@ 6.4. The dimension formula
The next formula is proved in [13]. It follows also from all our previous results.
Theorem 6.1.
$\sum_{0\leq i<\tau}\sigma(\nu_{P_{i}}(\Gamma_{i})+\delta_{i}),$
where $P_{0},\ldots,P_{\tau}$ are infinitely near points of $\Gamma$, $\tau$ is the first index such that $\Gamma_{\tau}$ is smooth, and $\delta_{i}$ is the number of irreducible components of $E_{i}$ passing through $P_{i}$.
Proof.
Notice that $\sigma(2)=\sigma(3)=0$, so that we do not need to assume that $\Gamma_{\tau}$ is transverse to $E_{\tau}$, just that it is non-singular. Up to replacing $\Gamma$ with another curve in its equisingularity class ${\mathcal{C}}$, we can assume that $\Gamma_{i}$ is generic in the equisingularity class of $(\Gamma_{i},E_{i})$ for any $0\leq i<\tau$ by Remark 5.8 . Now, Propositions 5.4 and 6.1 imply the result. ∎

### v3

@@ 6.2.3. $\delta_{y}=1$ and $\beta_{0}=n$
Again, we have $l_{0}=0$ for $\mathcal{C}$, $E=\{xy=0\}$ and ${E}_{1}=\{x_{1}=0\}$. As described in subsubsection 4.2.2, ${\mathcal{T}}_{0}=(\Omega_{1},\Omega_{2})$ where $\Omega_{1}=xdy\in\Delta_{n,<}^{0}$, $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{\mathcal{C}}(\Omega_{1})=2n$ and $\nu_{\mathcal{C}}(\Omega_{2})=n+\beta_{1}$. We also have $\tilde{\mathcal{T}}_{0}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$ where $\tilde{\Omega}_{1}=dx_{1}\in\Delta_{n,<}$, $\tilde{\Omega}_{2}=x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{1})=n$ and $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2})=\beta_{1}$. Once more, the families $\mathcal{T}_{0}$ and $\tilde{\mathcal{T}}_{0}$ are both terminal, and parallel. Equation (35) for $j\in\{1,2\}$ holds too.
@@ 6.2.4. $\delta_{y}=1$ and $\beta_{0}\not\in(n)$
Take $1\leq k\leq 2{\nu}_{l+1}$, and assume $\Omega_{k}^{1}\neq\Omega_{k}^{0}$. Then there exists $1\leq j<k$ such that $\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}$. Since
$\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}=[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0})]_{n},$
@@ 6.4. The dimension formula
Now, we prove Theorem 1.3.
Proof of Theorem 1.3.
Up to replacing $\Gamma$ with another curve in its equisingularity class $\mathcal{C}$}, we can assume that $\Gamma_{i}$ is generic in the equisingularity class of $(\Gamma_{i},E_{i})$ for any $0\leq i<\tau$ by Remark 5.8 . Now, Remark 5.11 and Propositions 5.4 and 6.1 imply the result. ∎